OpenStudy (anonymous):

Can anyone show me how to find the GCF of 72 and 78?

OpenStudy (anonymous):

Hello, thanks for responding to my question.

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Hello and thanks for responding

OpenStudy (anonymous):

\[72 = 2 \times 2 \times 2 \times 3 \times3\]\[78 = 2 \times 3 \times 13\]what factors do 72 and 78 have in common?

OpenStudy (anonymous):

2 and 3

OpenStudy (anonymous):

so GCF = 2 * 3 ~

OpenStudy (anonymous):

Wowwww, so its the numbers 2 and 3, not just one of the numbers?

OpenStudy (anonymous):

2 times 3 = 6 = gcf ~ 6 is the gcf

OpenStudy (anonymous):

So is 6 the GCF for both numbers

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now that was a little confusing!! lol

OpenStudy (anonymous):

72 can divide by 1, 2, 3, 4, 6, 8, 12, 18, 24, 36, 72 78 can divide by 1, 2, 3, 6, 13, 26, 39, 78. so 6 is the greatest common factors

OpenStudy (anonymous):

the way i did earlier was